# FlyLoRA: Boosting Task Decoupling and Parameter Efficiency via Implicit Rank-Wise Mixture-of-Experts

### v1

@@ 1 Introduction
However, to achieve strong performance on complex tasks, LoRA typically requires much higher ranks, which contradicts PEFT’s core goal of efficiency [31, 44]. Moreover, interference within LoRA’s ranks can impair training [75], leading to issues such as hallucination [22] and gradient explosion [57], thereby largely limiting its potential. We refer to this challenge as intra-task interference. Meanwhile, foundation models often need to integrate multiple capabilities to handle complex downstream tasks, but retraining on multi-domain corpora is expensive [70], especially since we already have several models that excels at specialized tasks. Consequently, model merging [11, 28, 29, 53] is widely used to combine LoRA components trained on different domains in a training-free manner. Arising from conflicts between different components, this process introduces another challenge: inter-task interference.
To address intra-task interference, several studies have incorporated the Mixture-of-Experts (MoE) architecture [19, 30, 34, 69, 77, 80] into LoRA [17, 20, 21, 37, 54, 75, 81, 88], where each expert learns specialized knowledge to partially achieve task decoupling. We refer to these approaches as MoE-based LoRA methods. They replace the original low-rank matrices with multiple experts and use a dynamic router to selectively activate them. By leveraging redundant parameters and sparse activations, these methods keep the computational budget comparable to LoRA with fewer total ranks. However, they may still suffer from interference within each expert. To investigate this issue, we conduct pivot studies on MMLU using Llama-3.1-8B, adopting Split-LoRA as a representative MoE-based method. As shown in Figure 1(a), finer-grained rank allocation yields consistent performance improvements under a fixed budget. However, as illustrated in Figure 1(b), pushing expert granularity to extremes also increases the number of activated trainable parameters due to additional router overhead. This trade-off makes it difficult to achieve both high performance and efficiency. Meanwhile, resolving inter-task interference has received limited emphasis in existing MoE-based LoRA methods.
@@
Inspired by the fly olfactory circuit [6, 13, 41, 71, 98], which shows strong similarity to MoE-based LoRA, we introduce an implicit router to mitigate the trade-off between intra-task interference and efficiency. As shown in Figure 1(c), this leads to the design of FlyLoRA, which (1) treats matrix $\bm{A}$ as a frozen sparse random projection that maps inputs into a higher-rank space (e.g., $r=32$ vs. $r=8$); and (2) simulates the bio-inspired “winner-take-all” mechanism by activating $k$ rank-1 experts in $\bm{B}$ linked to the top-$k$ magnitudes after projection by $\bm{A}$. This unifies the roles of $\bm{A}$ and router $\bm{G}$ into a single frozen projection, jointly performing down-projection and expert selection. Without explicit router parameters, the resulting implicit rank-wise MoE structure maintains computational efficiency while reducing intra-task interference. Moreover, we theoretically show that distinct random projections $\bm{A}_{i}$ and $\bm{A}_{j}$ from different LoRA components naturally map task updates into approximately orthogonal subspaces, thereby alleviating inter-task interference.
@@
  1. 1.
    Efficient Intra-task Decoupling: By using implicit rank-wise MoE, we enable finer expert allocation with reduced parameter interference in single-task scenarios. Additionally, FlyLoRA surpasses MoE-based LoRA in efficiency by eliminating the need for router parameters.
  2. 2.
    Efficient Inter-task Decoupling: In multi-task model merging scenarios, different random projections $\bm{A}_{i}$ and $\bm{A}_{j}$ naturally form approximately orthogonal subspaces. This inherent property ensures different LoRA components operate in uncorrelated subspaces, thus achieves decoupling.
  3. 3.
    Neuroscience-Inspired Design: The efficacy of our algorithm, combined with its structural alignment with the fly olfactory circuit, establishes a promising bridge between neuroscience and artificial intelligence.
@@ 2.2 MoE-based LoRA Framework
$f_{\text{MoE-LoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\sum_{i=1}^{N}\bm{G}(\bm{x})_{i}\cdot\underbrace{\bm{B}_{i}\bm{A}_{i}\bm{x}}_{\bm{E}_{i}(\bm{x})},$ (3)
@@ 2.3 Pushing MoE-based LoRA Architecture to the Extreme
$f_{\text{rank-wise-LoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\sum_{i=1}^{r}\bm{G}(\bm{x})_{i}\cdot\underbrace{\bm{b}_{i}\bm{a}_{i}\bm{x}}_{\bm{E}_{i}(\bm{x})},$ (5)
@@
$f_{\text{LoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\sum_{i=1}^{r}\bm{b}_{i}\bm{a}_{i}\bm{x}.$ (6)
Nevertheless, a naive approach of first computing all $r$ terms $\bm{b}_{i}\bm{a}_{i}\bm{x}$ and then selecting the top-$k$ would also forfeit the computational benefits of sparse activation, as the cost of computing all terms remains $\mathcal{O}(rmn)$. This necessitates a routing strategy that can identify the most important experts before fully computing their outputs. Furthermore, beyond efficient routing, another critical limitation of existing MoE-based LoRA methods is their lack of inherent support for multi-task deployment. When merging models already fine-tuned on different tasks, interference between experts often leads to significant performance degradation, as the underlying architecture does not structurally encourage task-specific updates to reside in orthogonal or non-overlapping parameter subspaces.
@@ 3 FlyLoRA
Inspired by the fly olfactory circuit (Figure 1(c)), whose neural architecture inherently meets our requirements for MoE-based LoRA variants, we propose FlyLoRA. Section 3.1 presents its formal design, while subsequent sections analyze its key advantages: Section 3.2 shows how a fixed $\bm{A}$ acts as an implicit router, Section 3.3 demonstrates intra-task decoupling, and Section 3.4 establishes inherent support for inter-task decoupling in model merging.
@@ 3.1 Formulation of FlyLoRA
$f_{\text{FlyLoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\Delta\bm{Wx}=\bm{W}_{0}\bm{x}+\sum_{i=1}^{r}\mathbb{I}(i\in\mathcal{I}_{\text{top}k})\cdot\bm{b}_{i}\bm{a}_{i}\bm{x},$ (11)
@@ 3.3 Gradient Decoupling via Top-k Sparsity
In the FlyLoRA framework, only matrix $\bm{B}$ requires updating. We theoretically demonstrate that our rank-wise expert allocation strategy, induced by top-$k$ selection, inherently reduces gradient covariance between distinct experts (see Theorem 3.2), thus mitigating intra-task interference.
@@ 3.4 Inter-Task Orthogonality in Model Merging
We analyze how FlyLoRA’s inherent subspace orthogonality enables effective multi-task model merging. Our theoretical foundation begins with Assumption 3.3 describing the sparsity pattern of activations. For analytical convenience, we consider a simplified condition where the top-$k$ operation randomly selects $k$ out of $r$ columns for activation.
Assumption 3.3 (Uniform Sparse Activation).
@@
For independent random matrices $\bm{A}_{i},\bm{A}_{j}\in\mathbb{R}^{r\times n}$ with sparse Gaussian entries ($\mathcal{N}(0,\frac{1}{r^{2}})$ for $p<n$ randomly selected entries per row), their column subspaces $\mathcal{S}_{i}=\text{span}(\bm{A}_{i})$ and $\mathcal{S}_{j}=\text{span}(\bm{A}_{j})$ satisfy:
@@
    Exponentially decaying correlations: $\mathbb{P}(\|\bm{A}_{i}\bm{A}_{j}^{\top}\|_{2}\geq\epsilon r)\leq\frac{p^{2}}{nr^{2}\epsilon^{2}}$
This theorem establishes that sparse random projections naturally induce nearly orthogonal subspaces. The residual correlation bound of order $\mathcal{O}(p^{2}/nr^{2})$ indicates that interference becomes negligible under practical parameter scales. This property directly leads to desirable behavior in model merging, as summarized in Corollary 3.5 (proof in Appendix A.3).
@@
This orthogonal decomposition provides a key advantage for model merging: task-specific updates $\bm{B}_{i}\bm{A}_{i}$ occupy nearly orthogonal subspaces, thereby preventing destructive interference when parameters are merged. The “Pairwise Orthogonality” property captures FlyLoRA’s fundamental behavior during multi-task aggregation. According to established geometric intuition that orthogonality facilitates model merging [29, 55, 74], and consistent with theoretical analyses in [38, 91], FlyLoRA’s fixed sparse projection design aligns with this principle. Empirically (see Section 4.3), the random projection $\bm{A}$ enables FlyLoRA to preserve task-specific performance after merging, whereas the learnable $\bm{A}$ in conventional LoRA exhibits significantly higher interference.
@@ 5.2 FlyLoRA’s Connection to Other Orthogonality-Based Designs in PEFT
Representative orthogonality-based PEFT methods like OFT [45, 58] and LoReFT [82] both operate on single tasks, and their orthogonal matrix $\bm{R}$ multiplies the pre-trained weight matrix $\bm{W}_{0}$, differing from LoRA variants (including FlyLoRA) that add $\Delta\bm{W}$ to $\bm{W}_{0}$. The multiplication scheme rotates the entire weight parameter space, and [45, 58] demonstrate this better adjusts semantic information compared to changing magnitude, explaining its success. In contrast, the additive scheme lacks this property since $\bm{W}_{0}$ cannot be rotated. In single-task settings, removing the MoE part and keeping only random $\bm{A}$ reduces FlyLoRA to LoRA-FA [93] or Asymmetry LoRA [96]. These variants can save resources but cannot improve performance. Thus, although all methods use orthogonality, FlyLoRA succeeds differently. We believe the orthogonality design in LoRA excels in multi-task scenarios, such as model merging (this work) and continual learning (O-LoRA [79]), because it decouples parameter interference across multiple downstream tasks when fine-tuning from the base model.
@@ 6 Related Work
The fly olfactory circuit [6, 13, 41, 71] serves as an exemplary model in bio-inspired AI due to its structural simplicity and functional completeness. Its core mechanism—random projection followed by sparse selection—effectively transforms high-dimensional inputs into separable representations. This biological principle has inspired algorithmic innovations across multiple AI domains, including locality-sensitive hashing [14, 66, 68], word embedding [40], federated learning [64], and continual learning [98]. The circuit’s enduring relevance highlights the value of cross-disciplinary inspiration in advancing computational methods.
@@ 8 Acknowledgments
We thank Cheems Wang for providing valuable suggestions on the manuscript.
@@ A.3 Random Projection Induces Approximate Subspace Orthogonality
For independent random matrices $\bm{A}_{i},\bm{A}_{j}\in\mathbb{R}^{r\times n}$ with sparse Gaussian entries ($\mathcal{N}(0,\frac{1}{r^{2}})$ for $p<n$ randomly selected entries per row), their column subspaces $\mathcal{S}_{i}=\text{span}(\bm{A}_{i})$ and $\mathcal{S}_{j}=\text{span}(\bm{A}_{j})$ satisfy:
@@
    Exponentially decaying correlations: $\mathbb{P}(\|\bm{A}_{i}\bm{A}_{j}^{\top}\|_{2}\geq\epsilon r)\leq\frac{p^{2}}{nr^{2}\epsilon^{2}}$
We demonstrate the approximate orthogonality between distinct LoRA components $\bm{B}_{i}\bm{A}_{i}$ and $\bm{B}_{j}\bm{A}_{j}$ through Frobenius inner product analysis:
@@
$\approx\sum_{i=1}^{t}w_{i}^{2}\|\bm{B}_{i}\bm{A}_{i}\|_{F}^{2}.$ (29)
We summarize these findings in the following corollary:
@@ Appendix D Limitations and Future Work
Recently, RL fine-tuning for LLMs has emerged as a promising approach that significantly enhances their reasoning ability [24]. However, stabilizing MoE RL training remains an open question [95], and further exploration will focus on the integration of FlyLoRA with RL training [60] and potentially extending it to offline policy optimization [49, 50, 51, 52, 59, 67]. Additionally, integrating active data selection methods could be a promising direction to further improve data efficiency [61, 76, 78].

### v2

@@ 1 Introduction
However, to achieve strong performance on complex tasks, LoRA typically requires much higher ranks, which contradicts PEFT’s core goal of efficiency [31, 45]. Moreover, interference within LoRA’s ranks can impair training [76], leading to issues such as hallucination [22] and gradient explosion [58], thereby largely limiting its potential. We refer to this challenge as intra-task interference. Meanwhile, foundation models often need to integrate multiple capabilities to handle complex downstream tasks, but retraining on multi-domain corpora is expensive [71], particularly when several specialized models are already available. Consequently, model merging [11, 28, 29, 54] is widely used to combine LoRA components trained on different domains in a training-free manner. Arising from conflicts between different components, this process introduces another challenge: inter-task interference.
To address intra-task interference, several studies have incorporated the Mixture-of-Experts (MoE) architecture [19, 30, 34, 70, 78, 81] into LoRA [17, 20, 21, 37, 55, 76, 82, 89], where each expert learns specialized knowledge to partially achieve task decoupling. We refer to these approaches as MoE-based LoRA methods. They replace the original low-rank matrices with multiple experts and use a dynamic router to selectively activate them. By leveraging redundant parameters and sparse activations, these methods keep the computational budget comparable to LoRA with fewer total ranks. However, they may still suffer from interference within each expert. To investigate this issue, we conduct pilot studies on MMLU using Llama-3.1-8B, adopting Split-LoRA as a representative MoE-based method. As shown in Figure 1(a), finer-grained rank allocation yields consistent performance improvements under a fixed budget. However, as illustrated in Figure 1(b), pushing expert granularity to extremes also increases the number of activated trainable parameters due to additional router overhead. This trade-off makes it difficult to achieve both high performance and efficiency. Meanwhile, resolving inter-task interference has received limited emphasis in existing MoE-based LoRA methods.
@@
Inspired by the fly olfactory circuit [6, 13, 38, 42, 72, 100], which shows strong similarity to MoE-based LoRA, we introduce an implicit router to mitigate the trade-off between intra-task interference and efficiency. As shown in Figure 1(c), this leads to the design of FlyLoRA, which (1) treats matrix $\bm{A}$ as a frozen sparse random projection that maps inputs into a higher-rank space (e.g., $r=32$ vs. $r=8$); and (2) simulates the bio-inspired “winner-take-all” mechanism by activating $k$ rank-1 experts in $\bm{B}$ linked to the top-$k$ magnitudes after projection by $\bm{A}$. This unifies the roles of $\bm{A}$ and router $\bm{G}$ into a single frozen projection, jointly performing down-projection and expert selection. Without explicit router parameters, the resulting implicit rank-wise MoE structure maintains computational efficiency while reducing intra-task interference. Moreover, we theoretically show that distinct random projections $\bm{A}_{i}$ and $\bm{A}_{j}$ from different LoRA components naturally map task updates into approximately orthogonal subspaces, thereby alleviating inter-task interference.
@@
$\bullet$ Efficient Intra-task Decoupling: By using implicit rank-wise MoE, we enable finer expert allocation with reduced parameter interference in single-task scenarios. Additionally, FlyLoRA surpasses MoE-based LoRA in efficiency by eliminating the need for router parameters.
$\bullet$ Efficient Inter-task Decoupling: In multi-task model merging scenarios, different random projections $\bm{A}_{i}$ and $\bm{A}_{j}$ naturally form approximately orthogonal subspaces. This inherent property ensures different LoRA components operate in uncorrelated subspaces, thus achieves decoupling.
$\bullet$ Neuroscience-Inspired Design: The efficacy of our algorithm, combined with its structural alignment with the fly olfactory circuit, establishes a promising bridge between neuroscience and artificial intelligence.
@@ 2.2 MoE-based LoRA Framework
$f_{\text{MoE-LoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\frac{\alpha}{r}\sum_{i=1}^{N}\bm{G}(\bm{x})_{i}\cdot\underbrace{\bm{B}_{i}\bm{A}_{i}\bm{x}}_{\bm{E}_{i}(\bm{x})},$ (3)
@@ 2.3 Pushing MoE-based LoRA Architecture to the Extreme
$f_{\text{rank-wise-LoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\frac{\alpha}{r}\sum_{i=1}^{r}\bm{G}(\bm{x})_{i}\cdot\underbrace{\bm{b}_{i}\bm{a}_{i}\bm{x}}_{\bm{E}_{i}(\bm{x})},$ (5)
@@
$f_{\text{LoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\frac{\alpha}{r}\sum_{i=1}^{r}\bm{b}_{i}\bm{a}_{i}\bm{x}.$ (6)
Nevertheless, a naive approach of first computing all $r$ terms $\bm{b}_{i}\bm{a}_{i}\bm{x}$ and then selecting the top-$k$ would also forfeit the computational benefits of sparse activation, as the cost of computing all terms remains $\mathcal{O}(rmn)$. This necessitates a routing strategy that can identify the most important experts before fully computing their outputs. Furthermore, beyond efficient routing, another critical limitation of existing MoE-based LoRA methods is their lack of inherent support for multi-task deployment. When merging models already fine-tuned on different tasks, interference between LoRA adapters often leads to significant performance degradation, as the underlying architecture does not structurally encourage task-specific updates to reside in orthogonal or non-overlapping parameter subspaces.
@@ 3 FlyLoRA
Inspired by the fly olfactory circuit (Figure 1(c)), whose neural architecture inherently meets our requirements for MoE-based LoRA variants, we propose FlyLoRA (visualized in Figure 2(c)). Section 3.1 presents its formal design, while subsequent sections analyze its key advantages: Section 3.2 shows how a fixed $\bm{A}$ acts as an implicit router, Section 3.3 demonstrates intra-task decoupling, and Section 3.4 establishes inherent support for inter-task decoupling in model merging.
@@ 3.1 Formulation of FlyLoRA
$f_{\text{FlyLoRA}}(\bm{x})=\bm{W}_{0}\bm{x}+\Delta\bm{Wx}=\bm{W}_{0}\bm{x}+\frac{\alpha}{r}\sum_{i=1}^{r}\mathbb{I}(i\in\mathcal{I}_{\text{top}k})\cdot\bm{b}_{i}\bm{a}_{i}\bm{x},$ (11)
@@ 3.3 Gradient Decoupling via Top-k Sparsity
In the FlyLoRA framework, only matrix $\bm{B}$ requires updating. We theoretically demonstrate that our rank-wise expert allocation strategy, induced by top-$k$ selection, inherently reduces gradient covariance between distinct experts, thus mitigating intra-task interference. Our analysis begins with Assumption 3.2 describing the sparsity pattern of activations. For analytical convenience, we consider a simplified condition where the top-$k$ operation randomly selects $k$ out of $r$ columns for activation.
Assumption 3.2 (Uniform Sparse Activation).
@@
$\mathbb{E}[\tilde{\bm{\Sigma}}_{(i,j)}]\approx\mathbb{E}[\bm{\Sigma}_{(i,j)}]\cdot\frac{k^{2}}{r^{2}},\quad\forall i\neq j.$
@@ 3.4 Inter-Task Orthogonality in Model Merging
We analyze how FlyLoRA’s inherent subspace orthogonality enables effective multi-task model merging. We derive Theorem 3.4 (proof in Appendix A.3).
@@
For independent random matrices $\bm{A}_{i},\bm{A}_{j}\in\mathbb{R}^{r\times n}$ with sparse Gaussian entries ($\mathcal{N}(0,\frac{1}{r^{2}})$ for $p<n$ randomly selected entries per row), the following holds,
@@
    Polynomially decaying correlations: $\mathbb{P}(\|\bm{A}_{i}\bm{A}_{j}^{\top}\|_{2}\geq\epsilon r)\leq\frac{p^{2}}{nr^{2}\epsilon^{2}}$
This theorem establishes that sparse random projections naturally induce nearly orthogonal subspaces. The residual correlation bound of order $\mathcal{O}(\frac{p^{2}}{nr^{2}})$ indicates that interference becomes negligible under practical parameter scales. This property directly leads to Corollary 3.5 (proof in Appendix A.3).
@@
This orthogonal decomposition provides a key advantage for model merging: task-specific updates $\bm{B}_{i}\bm{A}_{i}$ occupy nearly orthogonal subspaces, thereby preventing destructive interference. The “Pairwise Orthogonality” property captures FlyLoRA’s behavior during multi-task aggregation. According to geometric intuition that orthogonality facilitates model merging [29, 56, 75], and consistent with theoretical analyses in [39, 92], FlyLoRA’s fixed sparse projection design aligns with this principle. Empirically (see Section 4.3), the random projection $\bm{A}$ enables FlyLoRA to preserve task-specific performance after merging, whereas the learnable $\bm{A}$ in conventional LoRA exhibits significantly higher interference. A similar analysis of LoRA merging was conducted in a concurrent study [93].
@@ 5.2 FlyLoRA’s Connection to Other Orthogonality-Based Designs in PEFT
Representative orthogonality-based PEFT methods like OFT [46, 59] and LoReFT [83] both operate on single tasks, and their orthogonal matrix $\bm{R}$ multiplies the pre-trained weight matrix $\bm{W}_{0}$, differing from LoRA variants (including FlyLoRA) that add $\Delta\bm{W}$ to $\bm{W}_{0}$. The multiplication scheme rotates the entire weight parameter space, and [46, 59] demonstrate this better adjusts semantic information compared to changing magnitude, explaining its success. In contrast, the additive scheme lacks this property since $\bm{W}_{0}$ cannot be rotated. In single-task settings, removing the MoE part with only random $\bm{A}$ reduces FlyLoRA to LoRA-FA [94] or Asymmetry LoRA [97]. These variants can save resources but cannot improve performance. Thus, although all methods use orthogonality, FlyLoRA succeeds differently. We think the orthogonality design in LoRA excels in multi-task scenarios, such as model merging (this work and LoRI [93]) and continual learning (O-LoRA [80]), because it decouples parameter interference across multiple downstream tasks when fine-tuning from the base model.
@@ 6 Related Work
The fly olfactory circuit [6, 13, 38, 42, 72] serves as an exemplary model in bio-inspired AI due to its structural simplicity and functional completeness. Its core mechanism—random projection followed by sparse selection—effectively transforms high-dimensional inputs into separable representations. This biological principle has inspired algorithmic innovations across multiple AI domains, including locality-sensitive hashing [14, 67, 69], word embedding [41], federated learning [65], and continual learning [100, 101]. The circuit’s enduring relevance highlights the value of cross-disciplinary inspiration in advancing computational methods.
@@ 8 Acknowledgments
We thank Cheems Wang for his valuable suggestions on the manuscript. We also thank the anonymous reviewers for their positive feedback and constructive comments. This work was supported by the National Key R&D Program of China under Grant 2018AAA0102801.
@@ A.3 Random Projection Induces Approximate Subspace Orthogonality
For independent random matrices $\bm{A}_{i},\bm{A}_{j}\in\mathbb{R}^{r\times n}$ with sparse Gaussian entries ($\mathcal{N}(0,\frac{1}{r^{2}})$ for $p<n$ randomly selected entries per row), the following holds:
@@
    Polynomially decaying correlations: $\mathbb{P}(\|\bm{A}_{i}\bm{A}_{j}^{\top}\|_{2}\geq\epsilon r)\leq\frac{p^{2}}{nr^{2}\epsilon^{2}}$
We demonstrate the approximate orthogonality between distinct LoRA components $\bm{B}_{i}\bm{A}_{i}$ and $\bm{B}_{j}\bm{A}_{j}$ through Frobenius inner product analysis following [75]:
@@
$\approx\sum_{i=1}^{t}w_{i}^{2}\|\bm{B}_{i}\bm{A}_{i}\|_{F}^{2},$ (28)
which aligns with the “Weight disentanglement” property for task arithmetic [56]. We summarize these findings in the following corollary:
@@ Appendix D Limitations and Future Work
Recently, RL fine-tuning for LLMs has emerged as a promising approach that significantly enhances their reasoning ability [24]. However, stabilizing MoE RL training remains an open question [96], and further exploration will focus on the integration of FlyLoRA with RL training [61] and potentially extending it to offline policy optimization [50, 51, 52, 53, 60, 68]. Additionally, integrating active data selection methods could be a promising direction to further improve data efficiency [62, 77, 79, 99].